If 100 grams of NaOH is dissolved in water to give a solution with a volume of 250ml, what is the formality of the solution?


1
Expert's answer
2021-03-04T06:07:55-0500

Formality can be calculated as follows:

F = n / V = m / (Mr × V)

where F - formality, n - number of moles, V - volume, m - mass, Mr - molecular weight.

From here:

F(NaOH) = m(NaOH) / (Mr(NaOH) × Vsolution) = 100 g / (40 g/mol × 0.250 L) = 10 F


Answer: 10 F